OpenStudy (anonymous):

find the area of the part of paraboloid x=4y^2+4z^2 that lies inside the cylinder z^2+y^2=3 @Hero @ganeshie8

ganeshie8 (ganeshie8):

where are u stuck ?

ganeshie8 (ganeshie8):

just the axes were changed... other than that the problem is bit straightforward right ?

ganeshie8 (ganeshie8):

Surface area = \[ \iint \limits_R \sqrt{f_y ^2 + f_z^2 + 1} ~dy dz\]

ganeshie8 (ganeshie8):

take the partials and plugin

ganeshie8 (ganeshie8):

Surface area = \[ \iint \limits_R \sqrt{(8y)^2 + (8z)^2 + 1} ~dy dz\] \[ \iint \limits_R \sqrt{64(y^2+z^2) + 1} ~dy dz\]

ganeshie8 (ganeshie8):

next look at the cylinder to setup the bounds

ganeshie8 (ganeshie8):

z^2 + y^2 <= 3 gives a shodow of disk of radius sqrt(3) in yz plane

ganeshie8 (ganeshie8):

Surface area = \[ \iint \limits_R \sqrt{(8y)^2 + (8z)^2 + 1} ~dy dz\] \[ \iint \limits_R \sqrt{64(y^2+z^2) + 1} ~dy dz\] \[ \int \limits_0^{2\pi} \int \limits_0^{\sqrt{3}} \sqrt{64(r^2) + 1} ~r ~dr d\theta\]

ganeshie8 (ganeshie8):

evaluate

OpenStudy (anonymous):

how did you plug it in to wolffram aplha last time?

OpenStudy (anonymous):

thannks